User Tools

Site Tools


lbaops:lbafeb2008:todo

Parkes

  • Test Apple Xserve RAIDS DONE
  • Connect DAS to pkvsi1 (not v1pks)
  • Check VSIC connections
  • Test DAS

Mopra

  • Copy DAS profiles from v1mop DONE
  • Reconfig network setup DONE
  • Get turret rotation/translation working again
  • Verify 12mm Rx setup/operation
  • Verify all updated software working.
  • Setup hybrid for circular polarisation at 12mm. May have to be done on Wedn.

ATCA'

  • Reconnect VSIC and test DONE
  • Test dodgey channel on VSIC/cables DONE

Hobart

  • Check CVS version DONE

Ceduna

  • Check CVS version
  • Check can connect to green

General

  • Create VC fringes test experiments
  • Check CVS versio upto date
  • Merge UDP version?
  • Check cdisko working OK
  • Investigate cdisko scheduling
  • Check RtFC works with green
  • Copy schedules to all observatories and run atcasched at Mopra/ATCA
  • Label all disks
  • Update the disk status page on the wiki with disk IDs
lbaops/lbafeb2008/todo.txt · Last modified: 2015/12/18 16:38 (external edit)